OpenStudy (anonymous):

find polynomial function of degree 4 with -1 as zero of multiplicity 3 and 0 as a zero of multiplicity 1

OpenStudy (ipwnbunnies):

f(x) = (x+1)(x+1)(x+1)(x) I think that is it. Ya see, -1 can be a "zero" 3 times, which means it has a multiplicity of 3. And 0 has the multiplicity of 1. When you multiply that out, you'll get a fourth degree polynomial because x*x*x*x = x^4

OpenStudy (anonymous):

so the answer would come out to x^4 +3x^3+3x^2

OpenStudy (ipwnbunnies):

I'm not sure, but that looks mighty tasty.

OpenStudy (whpalmer4):

No, you're missing a bit on the expanded polynomial...

OpenStudy (whpalmer4):

Right approach, however — just need to expand it correctly.

OpenStudy (anonymous):

aww i was just missing a one!

OpenStudy (whpalmer4):

Uh, tell me what you think the answer is...

OpenStudy (anonymous):

x^4+3x^3+3x^2+x

OpenStudy (whpalmer4):

and it may be "just missing a one" to you, but that makes it a completely different polynomial and set of zeros! Instead of the zeros you were looking for, \[x^4+3x^3+3x^2=0\]has zeros at \(x=0\) (multiplicity 2) and \(x = -\frac{1}{2}(3\pm i\sqrt{3})\) Yes, your new answer is much improved!

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

I just learned about Polynomials, i didn't pay attention but gosh that looks hard.